1) A 123g sample of iron (lll) sulfate is in a container at STP. What volume is the container?
2) How many molecules of iron (lll) sulfate are in the container in Problem 1?

Grade:9

1 Answers

Vikas TU
7 years ago
First part →
moles of Fe2(SO4)3 => 123/152 = > 0.809 moles.
At STP ,
V = 0.809 * 22.4 = > 18.12 litre.
 
Second part ==> 
No. of molecules => moles x Na => 0.809 * 6.023 * 10^23 = > 4.872 x 10^23 molecules.